Home‎ > ‎Crew 158‎ > ‎

1211-XOReport

MISSION END

12/11/2015 / Fri

HSO/Philippe A. Souvestre

 

Second in Command: Nicholas Ip Jewell

Summary Title

 

MDRS CREW 158 has safely landed back on Earth

 

 

MISSION  STATUS: 

Overall, Mission status is evolving well as planned.

 

Morning briefings focused on 

1.       Transfer process regarding mission end, HAB command, and crew duties

2.       Ongoing pre-pilot exploratory biomedical feasibility evaluations

 

Mission planning for the day re

 

1.       0800 - Breakfast/Crew tasks and duties 

2.       Water system Logistics status review re crew showers, dish washing, loft tank maintenance

3.       EVA 10 of the day was cancelled due to no CAPCOM bandwidth available. It would have been led by Philippe Souvestre, with Nicholas Ip Jewell and Emmy Ip Jewell

4.       EVA replaced by cleaning and tiding HAB

5.       1300 – Lunch time and early afternoon were spent in Hanksville to ensure Wi-Fi capability to send all of yesterday and today reports

6.       1400 to 1700– Crew Reports writing

7.       1700 – Return to HAB

8.       1800 – 1845 – HAB Space arrangement and further cleaning

9.       1900 – 2100 – CAPCOM time window

10.    2100 – 2130 – Dinner time

11.    Crew group meditation/yoga and relaxation time

 

 

SIM HAB and Crew preparedness, activities, and actions went well as expected.

 

SOL  ACTIVITY   SUMMARY:

1.       HAB space and appliances review and tidiness

2.       Further implementation of three pre-pilot preliminary feasibility evaluations

 

Look Ahead Plan:

 

1.       How are you going to follow up? 

All studies, projects and objectives have been completed. Mission is accomplished.  

 

2.       What is the plan for tomorrow?

Tomorrow will be dedicated to proceeding for transfer of HAB command and duties to incoming Crew 159.

 

 

Anomalies and any problems in work:  

None

 

WeatherWet cold, no wind, complete low ceiling overcast skies. Ice and snow mixed throughout the whole day, no snow on the ground.

Crew Physical Status:  Everyone is in excellent health and still demonstrates first-rate morale at Mission End.


Reports to be filed:
Commander

Engineer

Journalist

EVA request

Daily Summary Report

Support Requested:    

None